# Enter the Mind Palace: Reasoning and Planning for Long-term Active Embodied Question Answering

### v1

@@ 4.1 Mind Palace Generation
The Robotic Mind Palace consists of a series of world instances representing the past long-term memory $[G_{1},\cdots,G_{N}]$ and the present knowledge of the environment $G_{0}$. At the start of the LA-EQA task, we assume the robot has not explored the present environment yet, so world instance $G_{0}$ is only initialized with area nodes $v$ because the state of the environment and object placement may have changed since the last mapping in $G_{1}$. We update $G_{0}$ as the robot explores the environment.
@@ 4.3 Early Stopping of Memory Retrieval for Navigation
This section examines how to reduce memory retrieval while maintaining exploration efficiency comparable to that of the unlimited memory retrieval case. In particular, we develop stopping criteria that decide when to halt past memory retrieval and proceed with exploration. Given a sequence of world instances that includes the present instance $G_{0}$ (e.g., $[G_{1},G_{2},G_{0}]$), we use an LLM to form a prediction set of areas $v\in G_{0}$, where the object $y$ can be located with a probability above a threshold $P(y)\geq 1-q$. Studies have shown that the LLM prediction and threshold $1-q$ can be calibrated [2, 50, 51]. Using the prediction set, we define two possible conditions to immediately stop memory retrieval from past world instances $[G_{1},G_{2}]$: 1) the prediction set contains only one area; 2) further memory retrieval will not improve the robot plan over the next sequence to explore $v_{i}$ in the prediction set. We evaluate the possible improvements on the sequence using the notion of Value of Information (VoI) [52], which quantifies the expected utility gain from retrieving past memory, reducing the expected exploration cost $J$. We discuss more details in Appendix B.
@@ 8 Limitations
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ Acknowledgments
Appendix A: Extended Problem Formulation of LA-EQA A
Appendix B: Mind Palace Exploration Details B
Appendix C: Long-term Active EQA Benchmark C
Appendix D: EQA Methods Details D
Appendix E: Evaluation Metrics E
Appendix F: Example Results from Different Methods F
Appendix G: Hardware Experiment Details G
Appendix H: Examples of Failure Cases H
Appendix I: Example of Full Results I
@@ B.3 Implementation Details Planning over episodic world instances
Prompt 4 (Continued) Will it reduce time and energy needed to explore the present? If the question involves the current state, consider both past and present. Preferred strategy order: PAST_ONLY >>> PAST_THEN_PRESENT >>> PRESENT_ONLY >>> MULTI_PAST_AND_PRESENT If selecting PAST_THEN_PRESENT or MULTI_PAST_AND_PRESENT, choose at most 5 past time instances. Based on your selected strategy, propose the ordered list of time instances to search that are most relevant. For context, here is the summary of the agent’s working memory and observations so far: {working_memory}
@@ B.4 Implementation Details Planning over Areas
We search for a sequence of areas $[v_{i},v_{i+1},\ldots]$ that minimizes the expected cost-to-go $J$. In our implementation, we perform a forward search up to three steps ahead to select the next area to explore, and return the first area $v_{i}$ in the optimal sequence.
@@ B.6 Early Stopping of Memory Retrieval for Navigation
We apply this principle to define early stopping conditions for searching for the object $y$ in the present environment $G_{0}$, beginning with memory retrieval from past world instances (e.g., $[G_{1},G_{2},G_{0}]$).
@@ B.7 Mind Palace Exploration algorithm
3:$\mathcal{M}\leftarrow\texttt{mind\_palace\_generation}(M,E)$
@@
16:$G\leftarrow\texttt{planning\_over\_world\_instances}(y,\mathcal{M},Q,h_{k})$
@@
29:    if $bool\_object\_found$:
@@
32:  if $bool\_object\_found$:

### v2

@@ 4.1 Mind Palace Generation
The Robotic Mind Palace consists of a series of world instances representing the past long-term memory $G_{1},\cdots,G_{N}]$ and the present knowledge of the environment $G_{0}$. At the start of the LA-EQA task, we assume the robot has not explored the present environment yet, so world instance $G_{0}$ is only initialized with area nodes $v$ because the state of the environment and object placement may have changed since the last mapping in $G_{1}$. We update $G_{0}$ as the robot explores the environment.
@@ 4.3 Early Stopping of Memory Retrieval for Navigation
This section examines how to reduce memory retrieval while maintaining exploration efficiency comparable to that of the unlimited memory retrieval case. In particular, we develop stopping criteria that decide when to halt past memory retrieval and proceed with exploration. Given a sequence of world instances that includes the present instance $G_{0}$ (e.g., $G_{1},G_{2},G_{0}]$), we use an LLM to form a prediction set of areas $v\in G_{0}$, where the object $y$ can be located with a probability above a threshold $P(y)\geq 1-q$. Studies have shown that the LLM prediction and threshold 1q1-q can be calibrated [2, 50, 51]. Using the prediction set, we define two possible conditions to immediately stop memory retrieval from past world instances $G_{1},G_{2}]$: 1) the prediction set contains only one area; 2) further memory retrieval will not improve the robot plan over the next sequence to explore $v_{i}$ in the prediction set. We evaluate the possible improvements on the sequence using the notion of Value of Information (VoI) [52], which quantifies the expected utility gain from retrieving past memory, reducing the expected exploration cost $J$. We discuss more details in Appendix B.
@@ 8 Limitations
Appendix A: Extended Problem Formulation of LA-EQA .A
Appendix B: Mind Palace Exploration Details .B
Appendix C: Long-term Active EQA Benchmark .C
Appendix D: EQA Methods Details .D
Appendix E: Evaluation Metrics .E
Appendix F: Example Results from Different Methods .F
Appendix G: Hardware Experiment Details .G
Appendix H: Examples of Failure Cases .H
Appendix I: Example of Full Results .I
@@ B.3 Implementation Details Planning over episodic world instances
Prompt 4 (Continued) Will it reduce time and energy needed to explore the present? If the question involves the current state, consider both past and present. Preferred strategy order: PAST_ONLY >> PAST_THEN_PRESENT >> PRESENT_ONLY >> MULTI_PAST_AND_PRESENT If selecting PAST_THEN_PRESENT or MULTI_PAST_AND_PRESENT, choose at most 5 past time instances. Based on your selected strategy, propose the ordered list of time instances to search that are most relevant. For context, here is the summary of the agent’s working memory and observations so far: {working_memory}
@@ B.4 Implementation Details Planning over Areas
We search for a sequence of areas $v_{i},v_{i+1},\ldots]$ that minimizes the expected cost-to-go $J$. In our implementation, we perform a forward search up to three steps ahead to select the next area to explore, and return the first area $v_{i}$ in the optimal sequence.
@@ B.6 Early Stopping of Memory Retrieval for Navigation
We apply this principle to define early stopping conditions for searching for the object $y$ in the present environment $G_{0}$, beginning with memory retrieval from past world instances (e.g., $G_{1},G_{2},G_{0}]$).
@@ B.7 Mind Palace Exploration algorithm
[…]
@@
[…]
@@
29:    if bool_object_foundbool\_object\_found:
@@
32:  if bool_object_foundbool\_object\_found: